OpenStudy (anonymous):

Find the surface area of the ball. Leave answers in terms of π. d =2 3/4 @Directrix

Directrix (directrix):

Surface Area of a Sphere - Answer in terms of PI A = 4 * π * r²

OpenStudy (anonymous):

so 4*pi*2 3/4?

Directrix (directrix):

diameter = 2 3/4 which is 11/4. half the diameter is 1/2 * 11/4 = 11/8

OpenStudy (anonymous):

this one is difficult to me since the 3/4 is there

Directrix (directrix):

Important Question: Is this --> d =2 3/4 2 and 3/4 OR is it 23 over 4. Very important to know. Look at the text.

OpenStudy (anonymous):

2 3/4

Directrix (directrix):

A = 4 * π * (11/8) ²

OpenStudy (anonymous):

this

Directrix (directrix):

Did you get ( 121π ) / 16 square units.

OpenStudy (anonymous):

i dont know how to put that into a calculator

Directrix (directrix):

It should be an answer option. Look and see. 121 pi all over 16

OpenStudy (anonymous):

oh yea it is

Directrix (directrix):

Good.